# A Robust Mixed-Effects Bandit Algorithm for Assessing Mobile Health Interventions

### v3

@@ D.1 Data from the IHS
We hereby define the reward, $R_{t}$, as the step count on the following day (cubic root). Our application also uses a subset of the baseline and contextual data; this subset contains the variables with the strongest association to the reward. Table 5 shows the features available to the bandit in the IHS data set.
@@
D.2 Evaluation
We run our contextual bandit on the IHS data, considering the binary action of whether or not to send a message at a given decision point based on the contextual variables $S_{1}$ and $S_{2}$. Each user is considered independently and with a cohesion network, for maximum personalization and independence of results. To guarantee that messages have a positive probability of being sent, we only sample the observations with notification randomization probability between $0.01$ and $0.99$. For the algorithm employing NNR, we defined participants in the same institution as their own “neighbors". This definition enables the flexibility for the value of $k$, representing the number of nearest neighbors, to vary for each participant based on their specific institutional context. Furthermore, in our study setting, we assume that individuals from the same "institution" enter the study simultaneously as a group. Due to the limited access to prior data, we are unable to build the neural linear models as in the Valentine Study.
We utilized 20 multiple-imputed data sets and performed an offline evaluation of the contextual bandit algorithms on each data set. The result is presented below in Figure 10. Similar to Section C.3, here we also compared DML-TS-NNR (as detailed in the main paper) with DML-TS-SU and NNR-Linear. The differences between these methods are illustrated in Figure. 11. In this specific dataset, both NNR and DML exhibit comparable performance individually, but their combined effect significantly enhances the overall performance of the algorithm.
Figure 10: (left) Unbiased estimates of the average per-trial reward for all four competing algorithms, relative to the reward obtained under the pre-specified Intern Health Study randomization policy across 20 multiple-imputed data sets. And (right) p-values from the pairwise paired t-tests. The dark shade in the last column indicates that the proposed DML-TS-NNR algorithm achieved significantly higher rewards than the other three competing algorithms while demonstrating comparable performance to the AC algorithm.
@@ D.2 Evaluation
Appendix E Additional Details for Algorithm 1
This appendix briefly discusses two details of Algorithm 1. The first is efficient computation of $V^{-1}$. Because $V$ is a fairly large matrix, a full matrix inversion is expensive. Fortunately, however, we can dramatically reduce the necessary computational requirements because each additional $(i,t)$ involves a rank-one perturbation to $V$. Consequently, we can apply the Sherman–Morrison formula to speed up computations. We leveraged this trick in our implementation of DML-TS-NNR for the simulation study by using efficient rank-one updates available in the SuiteSparse library (Davis & Hu, 2011).
@@ Appendix E Additional Details for Algorithm 1
$\mathcal{D}^{TS}$ is a multivariate distribution on $\mathbb{R}^{d}$ absolutely continuous with respect to Lebesgue measure which satisfies: 1. (anti-concentration) that there exists a strictly positive probability $p$ such that for any $u\in\mathbb{R}^{d}$ with $\|u\|=1$, $\mathbb{P}(u^{\top}\eta\geq 1)\geq p$; and 2. (concentration) there exists $c$, $c^{\prime}$ positive constants such that $\forall\delta\in(0,1)$, $P(\|\eta\|\leq\sqrt{cd\log(c^{\prime}d/\delta))}\geq 1-\delta$.
@@
Appendix F Regret Bound
F.1 Double Robustness of Pseudo-Reward
@@
$\mathbb{E}\left[\tilde{R}_{i,t}^{f}|s,\bar{a}\right]$ $=\Delta_{i,t}(s,\bar{a}).$
@@ F.2 Preliminaries
Let $X$ be a mean-zero sub-Gaussian random variable with variance factor $v^{2}$ and $Y$ be a bounded random variable such that $|Y|\leq B$ for some $0\leq B<\infty$. Then $XY$ is sub-Gaussian with variance factor $v^{2}B^{2}$.
@@
If $X,Y$ are sub-Gaussian with variance factors $v_{x}^{2}$, $v_{y}^{2}$, respectively, then $\alpha X+\beta Y$ is sub-Gaussian with variance factor $\alpha^{2}v_{x}^{2}+\beta^{2}v_{y}^{2}$$\forall\alpha,\beta\in\mathbb{R}$.
@@
Recall the equivalent definition of sub-Gaussianity that $X,Y$ are sub-Gaussian iff for some $b,a>0$ and all $\lambda>0$
$\mathbb{E}\exp\left(\lambda(X-\mathbb{E}X)\right)$ $\leq\exp(\lambda^{2}v_{x}^{2}/2)$
$\mathbb{E}\exp\left(\lambda(Y-\mathbb{E}Y)\right)$ $\leq\exp(\lambda^{2}v_{y}^{2}/2)$
@@
The following Lemma gives the sub-Gaussianity and variance of the difference between the pseudo-reward and its expectation. We see that in the variance, all terms except those involving the inverse propensity weighted noise variance vanish as $f_{i,t}$ becomes a better estimate of $r_{i,t}$. Note that means and variances may be implicitly conditioned on the history.
@@
If $\pi_{i,t}$ is correctly specified and $\tilde{\sigma}_{i,t}^{2}\geq\frac{1}{c}$, the difference between the pseudo-reward and its expectation (taken wrt the action and noise) is mean zero sub-Gaussian with variance
@@
Note that $|r_{i,t}(s,A_{i,t})|\leq\max\left(|r_{i,t}(s,\bar{a})|,|r_{i,t}(s,0)|\right)$ and $|f_{i,t}(s,A_{i,t})|\leq\max\left(|f_{i,t}(s,\bar{a})|,|f_{i,t}(s,0)|\right)$. Thus since $\left|\frac{1}{\delta_{A_{i,t}=\bar{a}}-\pi_{i,t}(0|s)}\right|$ is upper bounded by $c>0$, we have that $\frac{r_{i,t}(s,A_{i,t})-f_{i,t}(s,A_{i,t})}{\delta_{A_{i,t}=\bar{a}}-\pi_{i,t}(0|s)}$ is bounded and thus (not necessarily mean zero) sub-Gaussian. Since $\epsilon_{i,t}$ is sub-Gaussian, its denominator is bounded, and the remaining terms are deterministic, the entire difference between the pseudo-reward and its mean is sub-Gaussian. Now
@@
so that plugging this into Eqn. 14, we have
@@
and plugging this into Eqn. 13 we obtain the variance.
@@
and plugging this into Eqn. 15,
@@
$\mathbb{E}_{p(s,\bar{a})}\left[\textrm{Var}\left(\tilde{R}_{i,t}^{f}(s,\bar{a})\right)\right]$ $\leq\frac{\mathbb{E}_{p(s,\bar{a})}\left[(r_{i,t}(s,\bar{a})-f_{i,t}(s,\bar{a}))^{2}\right]\pi_{i,t}(0|s)+\textrm{Var}(\epsilon_{i,t})}{1-\pi_{i,t}(0|s)}$
$\qquad+\frac{\mathbb{E}_{p(s,\bar{a})}\left[(r_{i,t}(s,0)-f_{i,t}(s,0))^{2}\right][1-\pi_{i,t}(0|s)]+\textrm{Var}(\epsilon_{i,t})}{\pi_{i,t}(0|s)}$
$\qquad+2\sqrt{\mathbb{E}_{p(s,\bar{a})}\left[(r_{i,t}(s,\bar{a})-f_{i,t}(s,\bar{a}))^{2}\right]}\sqrt{\mathbb{E}_{p(s)}\left[(r_{i,t}(s,0)-f_{i,t}(s,0))^{2}\right]}$
@@
It follows immediately by taking expectations and applying Cauchy Schwartz. ∎
@@
$(r_{i,t}(\cdot,0)-f_{i,t}(\cdot,0))^{2}=o_{P}(k^{-1/2})$ and $(r_{i,t}(\cdot,\bar{a})-f_{i,t}(\cdot,\bar{a}))^{2}=o_{P}(k^{-1/2})$
@@
Note that
$P\left((r_{i,t}(\cdot,0)-f_{i,t}(\cdot,0))^{2}k^{1/2}>C\right)$ $\leq\frac{\mathbb{E}_{p(\mathcal{H}_{k-1})}\mathbb{E}_{p(s)}\left[(r_{i,t}(s,0)-f_{i,t}(s,0))^{2}\right]k^{1/2}}{C}$
@@
$=\frac{\mathbb{E}_{p(\mathcal{H}_{k-1})}o_{P}(1)}{C}$
$\leq\delta\textrm{ for sufficiently large $k$}$
so that $(r_{i,t}(\cdot,0)-f_{i,t}(\cdot,0))^{2}=o_{P}(k^{-1/2})$. In the second to last line we used that since $r_{i,t},f_{i,t}$ are bounded, $\mathbb{E}_{p(s)}\left[(r_{i,t}(s,0)-f_{i,t}(s,0))^{2}\right]$ is bounded. For bounded random variables, convergence in probability implies convergence in mean. A similar result holds when using $\bar{a}$ instead of 00. ∎
Corollary 3.
$\textrm{Var}\left(\tilde{R}_{i,t}^{f}(s,\bar{a})\right)$ $\leq\frac{o_{P}(k^{-1/2})\pi_{i,t}(0|s)+\textrm{Var}(\epsilon_{i,t})}{1-\pi_{i,t}(0|s)}+\frac{o_{P}(k^{-1/2})[1-\pi_{i,t}(0|s)]+\textrm{Var}(\epsilon_{i,t})+\textrm{Var}(\epsilon_{i,t})}{\pi_{i,t}(0|s)}$
$\qquad+2o_{P}(k^{-1/2})$
@@
This follows immediately from the previous two Corollaries. ∎
Corollary 4.
For all $\delta>0$, there exists $C>0$ s.t. w.p. at least $1-\delta$,
$\textrm{Var}\left(\tilde{R}_{i,t}^{f}(s,\bar{a})\right)$ $\leq\frac{Ck^{-1/2}\pi_{i,t}(0|s)+\textrm{Var}(\epsilon_{i,t})}{1-\pi_{i,t}(0|s)}+\frac{Ck^{-1/2}[1-\pi_{i,t}(0|s)]+\textrm{Var}(\epsilon_{i,t})}{\pi_{i,t}(0|s)}+2Ck^{-1/2}$
$\equiv v_{k}^{2}$
@@
By the definition of $o_{P}$, if $(f_{i,t}-r_{i,t})^{2}=o_{P}(k^{-1/2})$ for any $A>0,\delta>0$, there exists $K$ s.t. if $K>k$, then w.p. at least $1-\delta$,
$(f_{i,t}-r_{i,t})^{2}\leq Ak^{-1/2}.$
Then for all $k\in\mathbb{N}$, since $f,r$ are bounded
$(f_{i,t}-r_{i,t})^{2}$ $\leq 16B^{2}I(k\leq K)+I(k>K)Ak^{-1/2}.$
Now choose
$C$ $\equiv\max\left(A,16B^{2}K^{1/2}\right).$

### v4

@@ C.1 Data from the IHS
We define the reward, $R_{t}$, as the step count on the following day (cubic root). Our application also uses a subset of the baseline and contextual data; this subset contains the variables with the strongest association to the reward. Table 8 shows the features available to the bandit in the IHS data set.
@@
C.2 Evaluation
We run our contextual bandit on the IHS data, considering the binary action of whether or not to send a message at a given decision point based on the contextual variables $S_{1}$ and $S_{2}$. Each user is considered independently and with a cohesion network, for maximum personalization and independence of results. To guarantee that messages have a positive probability of being sent, we only sample the observations with notification randomization probability between $0.01$ and $0.99$. For the algorithm employing NNR, we defined participants in the same institution as their own “neighbors”. This definition enables the flexibility for the value of $k$, representing the number of nearest neighbors, to vary for each participant based on their specific institutional context. Furthermore, in our study setting, we assume that individuals from the same institution enter the study simultaneously as a group. Due to the limited access to prior data, we are unable to build the neural linear models as in the Valentine Study.
We utilized 20 multiple-imputed data sets and performed an offline evaluation of the contextual bandit algorithms on each data set. The result is presented below in Figure 11. Similar to Section B.4, here we also compared RoME (as detailed in the main paper) with RoME-SU and NNR-Linear. The differences between these methods are illustrated in Figure. 12. In this specific dataset, both NNR and DML exhibit comparable performance individually, but their combined effect significantly enhances the overall performance of the algorithm.
Figure 11: (left) Unbiased estimates of the average per-trial reward for all four competing algorithms, relative to the reward obtained under the pre-specified Intern Health Study randomization policy across 20 multiple-imputed data sets. And (right) p-values from the pairwise paired t-tests. The dark shade in the last column indicates that the proposed RoME algorithm achieved significantly higher rewards than the other three competing algorithms while demonstrating comparable performance to the AC algorithm.
@@ C.2 Evaluation
Appendix D Additional Details for Algorithm 1
This appendix briefly discusses two details of Algorithm 1. The first is efficient computation of $V^{-1}$. Because $V$ is a large matrix, a full matrix inversion is expensive. Fortunately, however, we can dramatically reduce the necessary computational requirements because each additional $(i,t)$ involves a rank-one perturbation to $V$. Consequently, we can apply the Sherman–Morrison formula to speed up computations. We leveraged this trick in our implementation of RoME for the simulation study by using efficient rank-one updates available in the SuiteSparse library (Davis & Hu, 2011).
@@ Appendix D Additional Details for Algorithm 1
$\mathcal{D}^{TS}$ is a multivariate distribution on $\mathbb{R}^{d}$ absolutely continuous with respect to Lebesgue measure which satisfies: 1. (anti-concentration) that there exists a strictly positive probability $p$ such that for any $u\in\mathbb{R}^{d}$ with $\|u\|=1$, $\mathbb{P}(u^{\top}\eta\geq 1)\geq p$; 2. (concentration) there exists $c$, $c^{\prime}$ positive constants such that $\forall\delta\in(0,1)$, $P(\|\eta\|\leq\sqrt{cd\log(c^{\prime}d/\delta))}\geq 1-\delta$; and 3. it possesses a finite second moment, $\mathbb{E}\|\eta\|^{2}$, where $\eta\sim\mathcal{D}^{TS}$.
@@
Appendix E Regret Bound
E.1 Double Robustness of Pseudo-Reward
Going forward, we use the notation $\Delta^{f}_{i,t}(s,\bar{a})\coloneq f_{i,t}(s,\bar{a})-f_{i,t}(s,0)$ to denote the prediction of the differential reward. In the computations in this section, we implicitly condition on $\mathcal{H}_{i,t}$ and $A_{i,t}\in\{0,\bar{a}\}$.
@@ E.1 Double Robustness of Pseudo-Reward
$\mathbb{E}\left(\tilde{R}_{i,t}^{f}|s,\bar{a}\right)$ $=\Delta_{i,t}(s,\bar{a}).$
@@ E.2 Preliminaries
Let $X$ be a mean-zero sub-Gaussian random variable with variance proxy $v^{2}$ and $Y$ be a bounded random variable such that $|Y|\leq B$ for some $0\leq B<\infty$. Then $XY$ is sub-Gaussian with variance proxy $v^{2}B^{2}$.
@@
If $X,Y$ are sub-Gaussian with variance proxies $v_{x}^{2}$, $v_{y}^{2}$, respectively, then $\alpha X+\beta Y$ is sub-Gaussian with variance proxy $\alpha^{2}v_{x}^{2}+\beta^{2}v_{y}^{2}$$\forall\alpha,\beta\in\mathbb{R}$.
@@
Recall the equivalent definition of sub-Gaussianity that $X,Y$ are sub-Gaussian iff for some $a,b>0$ and all $\lambda>0$
$\mathbb{E}\exp\left(\lambda(X-\mathbb{E}X)\right)$ $\leq\exp(\lambda^{2}v_{x}^{2}/2),$
$\mathbb{E}\exp\left(\lambda(Y-\mathbb{E}Y)\right)$ $\leq\exp(\lambda^{2}v_{y}^{2}/2).$
@@
The following Lemma gives the sub-Gaussianity and variance of the difference between the pseudo-reward and its expectation. We see that in the variance, all terms except those involving the inverse propensity weighted noise variance vanish as $f_{i,t}$ becomes a better estimate of $r_{i,t}$.
@@
The difference between the pseudo-reward and its expectation (taken wrt the action and noise) is mean zero sub-Gaussian with variance
@@
Note that $|r_{i,t}(s,A_{i,t})|\leq\max\left(|r_{i,t}(s,\bar{a})|,|r_{i,t}(s,0)|\right)\leq M$ and $|f_{i,t}(s,A_{i,t})|\leq\max\left(|f_{i,t}(s,\bar{a})|,|f_{i,t}(s,0)|\right)\leq M$. Thus, since $\left|1/\{\delta_{A_{i,t}=\bar{a}}-\pi_{i,t}(0|s)\}\right|$ is upper bounded (because $\pi_{i,t}(0|s)\in[\pi_{\min},\pi_{\max}]$), we have that $\frac{r_{i,t}(s,A_{i,t})-f_{i,t}(s,A_{i,t})}{\delta_{A_{i,t}=\bar{a}}-\pi_{i,t}(0|s)}$ is bounded and thus (not necessarily mean zero) sub-Gaussian. Since $\epsilon_{i,t}$ is sub-Gaussian, its denominator is bounded, and the remaining terms are deterministic, the entire difference between the pseudo-reward and its mean is sub-Gaussian. Now
@@
so that plugging this into (8), we have
@@
and plugging this into (7) we obtain the variance.
@@
and plugging this into (9),
@@
Let $\tilde{\pi}\coloneq\min(\pi_{\min},1-\pi_{\max})$ and $\sigma^{2}\coloneq\textrm{Var}(\epsilon_{i,t})$. Then
$\textrm{Var}\left(\tilde{R}_{i,t}^{f}(s,\bar{a})\right)\leq\frac{2\sigma^{2}+4M^{2}}{\tilde{\pi}}+8M^{2}\eqcolon v_{1}^{2}.$
@@
Substitute $\sigma^{2}\coloneq\textrm{Var}(\epsilon_{i,t})$ in Lemma 5 and apply the bounds $M$, $\pi_{\min}$, and $\pi_{\max}$ to obtain
$\textrm{Var}\left(\tilde{R}_{i,t}^{f}(s,\bar{a})\right)\leq$ $\frac{4M^{2}\pi_{i,t}(0|s)+\sigma^{2}}{\tilde{\pi}}+\frac{4M^{2}\{1-\pi_{i,t}(0|s)\}+\sigma^{2}}{\tilde{\pi}}+8M^{2}$
=\ = $\frac{2\sigma^{2}+4M^{2}}{\tilde{\pi}}+8M^{2}.$
We now show that Assumption 3 results in a lower asymptotic variance bound.
@@
The asymptotic variance of $\tilde{R}_{i,t}^{f}(s,\bar{a})$ is no greater than
$v_{2}^{2}\coloneq\frac{\sigma^{2}}{\tilde{\pi}(1-\tilde{\pi})}.$
@@
By Assumption 3, the terms involving $r_{i,t}(s,\bar{a})-f_{i,t}(s,\bar{a})$ and $r_{i,t}(s,0)-f_{i,t}(s,0)$ converge almost surely to zero. Dropping these terms from the expression in Lemma 5, the asymptotic variance is then less than or equal to
$\frac{\sigma^{2}}{1-\pi_{i,t}(0|s)}+\frac{\sigma^{2}}{\pi_{i,t}(0|s)}=\frac{\sigma^{2}}{\pi_{i,t}(0|s)\{1-\pi_{i,t}(0|s)\}}\leq\frac{\sigma^{2}}{\tilde{\pi}(1-\tilde{\pi})}.$
Remark 1.
Comparing $v_{1}^{2}$ and $v_{2}^{2}$, we see that $v_{1}^{2}\geq v_{2}^{2}$ if
$\frac{2\sigma^{2}}{\tilde{\pi}}\geq\frac{\sigma^{2}}{\tilde{\pi}(1-\tilde{\pi})}\iff 1-\tilde{\pi}\geq 1/2\iff 1/2\geq\tilde{\pi},$
which must hold by the definition of $\tilde{\pi}$. Consequently, The effect of the DML is to lower the variance bound.
E.3 Derivation of Confidence Sets
To facilitate the developments below, we define the following quantities:
  • $V_{0,k},V_{i,t},b_{i,t},\hat{\theta}_{i,t}$: The quantities $V_{0}$, $V$, $b$, $\hat{\theta}\coloneq V^{-1}b$ in Algorithms 1 and 2 at decision point $(i,t)$.
  • $\mathcal{O}_{i,t}$: The set of user–time pairs up to but excluding $(i,t)$; we can express $\mathcal{O}_{i,t}$ as $\{(1,1),(1,2),(2,1),\ldots,(i,t)\}\backslash\{(i,t)\}$
  • $\theta_{i,t}^{\star}\coloneq\theta^{\text{shared}}+\theta^{\text{user}}_{i}+\theta^{\text{time}}_{t}=C_{i,t}\theta_{k}^{\star}\in\mathbb{R}^{d}$.
  • $\check{\theta}_{i,t}\coloneq C_{i,t}\hat{\theta}_{i,t}\in\mathbb{R}^{d}$.
  • $\Lambda_{i,t}^{0}\coloneq C_{i,t}V_{i,t}^{-1}V_{0,k}V_{i,t}^{-1}C_{i,t}^{{\top}}\in\mathbb{R}^{(d\times d)}$.
  • $\Lambda_{i,t}^{+}\coloneq C_{i,t}V_{i,t}^{-1}\left[\sum_{(j,u)\in\mathcal{O}_{i,t}}\tilde{\sigma}_{j,u}^{2}\phi\{x(S_{ju},A_{ju})\}\phi\{x(S_{ju},A_{ju})\}^{{\top}}\right]V_{i,t}^{-1}C_{i,t}^{{\top}}\in\mathbb{R}^{(d\times d)}$.
@@ E.3 Derivation of Confidence Sets
In the proof below, we allow the dimension of $\theta_{k}^{\star}$ to increase as stages progress so that $\theta_{k}^{\star},b_{i,t},\hat{\theta}_{i,t}\in\mathbb{R}^{(2k+1)d}$ and $V_{0,k},V_{i,t}\in\mathbb{R}^{(2k+1)d\times(2k+1)d}$. However, the proof is also applicable (with minor modifications) to settings where the final dimension of $\theta_{k}^{\star}$ is known and set in advance (as in Algorithm 2).
[…]
Assumption 10.
Let $\mathcal{O}$ be a set of $(i,t)$ pairs containing at least $K^{\star}\in\mathbb{N}$ elements, and let $\omega>0$. Then the following ordering holds:
$\sum_{(i,t)\in\mathcal{O}}x_{i,t}x_{i,t}^{{\top}}\succ(\#\mathcal{O})\omega I_{d},$
where #$\#\mathcal{O}$ denotes the cardinality of $\mathcal{O}$.
This assumption ensures that we continue acquiring information about all elements of $x_{i,t}$ at a linear rate as we progress through stages. We now adapt the classical theory on RLS estimation error to our setting.
Lemma 6 (Adapted from Theorem 2 in Abbasi-Yadkori et al. (2011)).
Let $\check{\theta}_{i,t}$ be the regularized least squares (RLS) estimate from Algorithm 1 for the $i$-th user at the $t$-th time point, corresponding to stage $k\coloneq i+t-1$ and let $\theta_{i,t}^{\star}$ be the true parameter value. For any $\delta>0$, with probability at least $1-\delta$ the estimates $\{\check{\theta}_{i,t}\}_{(i,t)\in\mathcal{O}_{K}}$ satisfy
$\|\check{\theta}_{i,t}-\theta_{i,t}^{\star}\|_{\underline{\hbox{\set@color$V$}}_{i,t}}\leq\ v\sqrt{2\log\left\{\frac{1}{\delta}\cdot\sqrt{\frac{\det(\underline{\hbox{\set@color$V$}}_{i,t}^{-1})}{\det(\Lambda_{i,t}^{0})}}\right\}}+\|C_{i,t}V_{i,t}^{-1}V_{0,k}\theta_{k}^{\star}\|_{\underline{\hbox{\set@color$V$}}_{i,t}},$
where $v^{2}$ is the variance proxy for the difference between the pseudo-reward and its mean.
@@
Let $m_{i,t}=\tilde{\sigma}_{i,t}\phi\{x(S_{i,t},A_{i,t})\}$ and $\rho_{i,t}=\tilde{\sigma}_{i,t}\{\tilde{R}_{i,t}^{f}(S_{i,t},\bar{A}_{i,t})-\mathbb{E}(\tilde{R}_{i,t}^{f}|S_{i,t},\bar{A}_{i,t})\}$. Further let
$\xi_{i,t}\coloneq$ :- $\sum_{(j,u)\in\mathcal{O}_{i,t}}m_{j,u}\rho_{j,u}\in\mathbb{R}^{(2k+1)d}.$
$\hat{\theta}_{i,t}$ $=V_{i,t}^{-1}b_{i,t}$
$=V_{i,t}^{-1}(\xi_{i,t}+\Phi_{i,t}^{\top}W_{i,t}\Delta_{i,t})$
$=V_{i,t}^{-1}\xi_{i,t}+V_{i,t}^{-1}\Phi_{i,t}^{{\top}}W_{i,t}\Phi_{i,t}\theta_{k}^{\star}$
$=V_{i,t}^{-1}\xi_{i,t}+V_{i,t}^{-1}(\Phi_{i,t}^{{\top}}W_{i,t}\Phi_{i,t}+V_{0,k})\theta_{k}^{\star}-V_{i,t}^{-1}V_{0,k}\theta_{k}^{\star}$
$=V_{i,t}^{-1}\xi_{i,t}+\theta_{k}^{\star}-V_{i,t}^{-1}V_{0,k}\theta_{k}^{\star}$
and thus
$\hat{\theta}_{i,t}-\theta_{k}^{\star}$ $=V_{i,t}^{-1}(\xi_{i,t}-V_{0,k}\theta_{k}^{\star}).$
Premultiplying, we can then obtain
$(\hat{\theta}_{i,t}-\theta_{k}^{\star})^{{\top}}C_{i,t}^{{\top}}\underline{\hbox{\set@color$V$}}_{i,t}C_{i,t}(\hat{\theta}_{i,t}-\theta_{k}^{\star})=(\hat{\theta}_{i,t}-\theta_{k}^{\star})^{{\top}}C_{i,t}^{{\top}}\underline{\hbox{\set@color$V$}}_{i,t}C_{i,t}V_{i,t}^{-1}(\xi_{i,t}-V_{0,k}\theta_{k}^{\star}).$
Simplification yields
$\|\check{\theta}_{i,t}-\theta_{i,t}^{\star}\|_{\underline{\hbox{\set@color$V$}}_{i,t}}^{2}=(\check{\theta}_{i,t}-\theta_{i,t}^{\star})^{{\top}}\underline{\hbox{\set@color$V$}}_{i,t}C_{i,t}V_{i,t}^{-1}(\xi_{i,t}-V_{0,k}\theta_{k}^{\star}).$
We can then apply the Cauchy-Schwarz inequality to bound the right-hand side.
$\|\check{\theta}_{i,t}-\theta_{i,t}^{\star}\|_{\underline{\hbox{\set@color$V$}}_{i,t}}^{2}\leq\|\check{\theta}_{i,t}-\theta_{i,t}^{\star}\|_{\underline{\hbox{\set@color$V$}}_{i,t}}\cdot\|C_{i,t}V_{i,t}^{-1}(\xi_{i,t}-V_{0,k}\theta_{k}^{\star})\|_{\underline{\hbox{\set@color$V$}}_{i,t}}$
Dividing by $\|\check{\theta}_{i,t}-\theta_{i,t}^{\star}\|_{\underline{\hbox{\set@color$V$}}_{i,t}}$ and applying the triangle inequality gives
$\|\check{\theta}_{i,t}-\theta_{i,t}^{\star}\|_{\underline{\hbox{\set@color$V$}}_{i,t}}\leq\|C_{i,t}V_{i,t}^{-1}\xi_{i,t}\|_{\underline{\hbox{\set@color$V$}}_{i,t}}+\|C_{i,t}V_{i,t}^{-1}V_{0,k}\theta_{k}^{\star}\|_{\underline{\hbox{\set@color$V$}}_{i,t}}$
We can show that
$(\Lambda_{i,t}^{0}+\Lambda_{i,t}^{+})^{-1}$
=\ = $\left\{C_{i,t}V_{i,t}^{-1}V_{0,k}V_{i,t}^{-1}C_{i,t}^{{\top}}+\sum_{(j,u)\in\mathcal{O}_{i,t}}C_{i,t}V_{i,t}^{-1}m_{j,u}m_{j,u}^{{\top}}V_{i,t}^{-1}C_{i,t}^{{\top}}\right\}^{-1}$
=\ = $\left\{C_{i,t}V_{i,t}^{-1}\left(V_{0,k}+\sum_{(j,u)\in\mathcal{O}_{i,t}}\tilde{\sigma}_{j,u}^{2}\phi_{j,u}\phi_{j,u}^{{\top}}\right)V_{i,t}^{-1}C_{i,t}^{{\top}}\right\}^{-1}$
=\ = $\left(C_{i,t}V_{i,t}^{-1}V_{i,t}V_{i,t}^{-1}C_{i,t}^{{\top}}\right)^{-1}$
=\ = $\left(C_{i,t}V_{i,t}^{-1}C_{i,t}^{{\top}}\right)^{-1}$
[…]
Combined with the fact that $\rho_{i,t}$ is sub-Gaussian with variance proxy $v^{2}$, the above result allows us to apply Theorem 1 of Abbasi-Yadkori et al. (2011). Specifically, to match their notation (setting aside our definition of $S_{i,t}$ for the moment), let $S_{i,t}=C_{i,t}V_{i,t}^{-1}\xi_{i,t}=\sum_{(j,u)\in\mathcal{O}_{i,t}}C_{i,t}V_{i,t}^{-1}\rho_{j,u}m_{j,u}$ and $\bar{V}_{i,t}=C_{i,t}V_{i,t}^{-1}V_{0,k}V_{i,t}^{-1}C_{i,t}^{\top}+C_{i,t}V_{i,t}^{-1}\left(\sum_{(j,u)\in\mathcal{O}_{i,t}}m_{j,u}m_{j,u}^{\top}\right)V_{i,t}^{-1}C_{i,t}^{\top}$, we have
$\|S_{i,t}\|^{2}_{\bar{V}_{i,t}}$ $\leq 2v^{2}\log\left(\frac{\det(\underline{\hbox{\set@color$V$}}_{i,t}^{-1})^{1/2}}{\delta\det(\Lambda_{i,t}^{0})^{1/2}}\right)$
$\|C_{i,t}V_{i,t}^{-1}\xi_{i,t}\|_{\underline{\hbox{\set@color$V$}}_{i,t}}$ $\leq v\sqrt{2\log\left\{\frac{1}{\delta}\cdot\sqrt{\frac{\det(\underline{\hbox{\set@color$V$}}_{i,t}^{-1})}{\det(\Lambda^{0}_{i,t})}}\right\}}$
Adding this bound to $\|C_{i,t}V_{i,t}^{-1}V_{0,k}\theta_{k}^{\star}\|_{\underline{\hbox{\set@color$V$}}_{i,t}}$ completes the proof.
[…]
Definition 1.
[…]
Lemma 7.
$\|C_{i,t}V_{i,t}^{-1}V_{0,k}\theta_{k}^{\star}\|_{\underline{\hbox{\set@color$V$}}_{i,t}}$ $\leq\|V_{0,k}\theta_{k}^{\star}\|_{V_{i,t}^{-1}}.$
@@
[…]
$V_{i,t}=\begin{bmatrix}Z_{11}&Z_{12}\\ Z_{21}&Z_{22}.\end{bmatrix}$
Using the $2\times 2$ block-matrix inversion identity, it then follows that
$V_{i,t}^{-1}=\begin{bmatrix}P^{-1}&-P^{-1}Z_{12}Z_{22}^{-1}\\ -Z_{22}^{-1}Z_{21}P^{-1}&Z_{22}^{-1}+Z_{22}^{-1}Z_{21}P^{-1}Z_{12}Z_{22}^{-1}\\ \end{bmatrix},$
where $P\coloneq Z_{11}-Z_{12}Z_{22}^{-1}Z_{21}$.
[…]
$\|C_{i,t}V_{i,t}^{-1}V_{0,k}\theta_{k}^{\star}\|^{2}_{\underline{\hbox{\set@color$V$}}_{i,t}}$
=\ = $\theta_{k}^{\star{\top}}V_{0,k}G^{{\top}}P^{-1}(1_{3}\otimes I_{d})\{(1_{3}^{{\top}}\otimes I_{d})P^{-1}(1_{3}\otimes I_{d})\}^{-1}(1_{3}^{{\top}}\otimes I_{d})P^{-1}GV_{0,k}\theta_{k}^{\star}$
[…]
where we have used the fact that $P^{-1/2}(1_{3}\otimes I_{d})\{(1_{3}^{{\top}}\otimes I_{d})P^{-1}(1_{3}\otimes I_{d})\}^{-1}(1_{3}^{{\top}}\otimes I_{d})P^{-1/2}\preceq I_{d}$ because the former is a projection matrix and thus its maximum eigenvalue is one. We can further simplify and bound this quantity as follows:
$\theta_{k}^{\star{\top}}V_{0,k}G^{{\top}}P^{-1}GV_{0,k}\theta_{k}^{\star}$
=\ = $\theta_{k}^{\star{\top}}V_{0,k}\begin{bmatrix}I_{3d}\\ -Z_{22}^{-1}Z_{21}\end{bmatrix}P^{-1}\begin{bmatrix}I_{3d}&-Z_{12}Z_{22}^{-1}\end{bmatrix}V_{0,k}\theta_{k}^{\star}$
=\ = $\theta_{k}^{\star{\top}}V_{0,k}\begin{bmatrix}P^{-1}&-P^{-1}Z_{12}Z_{22}^{-1}\\ Z_{22}^{-1}Z_{21}P^{-1}&Z_{22}^{-1}Z_{21}P^{-1}Z_{12}Z_{22}^{-1}\end{bmatrix}V_{0,k}\theta_{k}^{\star}$
=\ = $\theta_{k}^{\star{\top}}V_{0,k}\left(V_{i,t}^{-1}-\begin{bmatrix}0&0\\ 0&Z_{22}^{-1}\end{bmatrix}\right)V_{0,k}\theta_{k}^{\star}$
[…]